# Silence Speaks Volumes: A New Paradigm for Covert Communication via History Timing Patterns

### v1

@@ 5 Evaluation
Because analyzing large network captures is demanding in terms of computing resources and live experiments can only be conducted one after the other, we used genetic algorithms to find useful parameter combinations [genai24]. The Simple Genetic Algorithm (SGA) iteratively optimizes the selection of combinations of experimental parameters based on a performance metric computed from previous experiment runs called fitness. Each individual represents a distinct configuration of experimental parameters. At each iteration, the algorithm first selects a subset of elite individuals with the highest fitness scores. These elite configurations serve as parents in the crossover phase, where each offspring is generated by randomly inheriting parameter values from one of the two parents. To maintain diversity and explore new regions in the parameter space, a mutation operator is then applied to each offspring with a predefined mutation rate, potentially altering one or more parameter values. This straightforward evolutionary cycle—comprising selection, crossover, and mutation—enables the algorithm to progressively refine and schedule more promising parameter sets for subsequent experiments, while also ensuring that previously tested configurations are not redundantly re-evaluated.
@@
Live measurements were taken during the day and at night in a home network with 10 devices such as smartphones, smart home devices, and file servers. For offline analysis, traffic in this network was collected during the day and at night. In the WAN scenario, we used traffic recordings from the WIDE project [WIDE].
@@ 5.1 Environmental robustness
Robustness is crucial for covert channels to ensure reliable communication in real-world network environments where such impairments are common. This is especially true for timing-based channels, as network conditions affect packet timing and order [wendzelpatternsurvey15]. We evaluated the robustness of SHP under various network conditions, specifically focusing on the impact of delay, jitter, noise, and packet loss on bandwidth.
@@ 5.1.1 Impact of delay
The observed values for one-way network delay are typically in the ranges $\leq 110$ ms in wired LANs [delay14] and 5050300300 ms in WAN communication [delay12], depending on distance, routing, and network congestion. When using timing-based input sources, this causes differences in observed inputs between the CS and CR, especially when absolute timing is used as a reference for pointers.
In the WAN scenario, consistent network delay has a reduced impact on the robustness of SHP, as it can use relative timing (like in ISDs) or packet ordering (like in ISPNs) algorithms. This is because the CR computes intra-signal delays relative to previous pointers or initial synchronization. Let $ISD_{CS}(n)$ be the difference between when the CS observes a signal $n$ and the previous signal n1n-1, and the network delay between covert agents be $d$. Then the observed ISDs for both parties are $ISD_{CS}(n)=(t_{n}+d)-(t_{n-1}+d)=t_{n}-t_{n-1}+(d-d)=t_{n}-t_{n-1}=ISD_{CR}(n)$. This means that the observed interpacket delays or relative times remain the same, rendering any consistent network delay inconsequential for matching in the WAN scenario. This aligns with recent covert communication work in continuous or time-uncertain settings, where robustness is also achieved by leveraging timing uncertainty rather than attempting to tightly synchronize clocks [Lu2023TimeUncertainty, Li2022Jammer]. Similarly, for the ISPN algorithm, which relies on packet ordering, a constant delay does not alter the sequence of packet arrivals, so SHP remains robust under high network delays as long as they are stable.
@@ 5.1.2 Impact of jitter
We also implemented a silent interval using the method suggested by Wendzel et al. [DYST22]: If the CS receives two or more POI in less than $\phi$ milliseconds, all overt traffic in that time frame will be ignored, so that only isolated POI will be considered for matching. Consequently, the CR uses $\phi$ to ignore all POI received less than this number of milliseconds before a pointer and only interpret it as pointing to POI before that. This ensures the CS can calculate and send a pointer without running into a race condition with consecutive POI. Nonetheless, the larger we choose $\phi$, the more this will reduce matching chances and thus potential bandwidth.
@@
We tested the impact of jitter using simulations as well as live network channels. In the simulations, we produced challenging network circumstances for the LAN and WAN scenarios, using a rather high time difference between CS and CR of 2ms for the LAN scenario [thombre20] and 20ms for the WAN scenario [jitter11] and only measured the impact of rounding network jitter of overt traffic. The results show that while less precise rounding improves performance when the jitter is negligible, more aggressive rounding improves performance when the jitter is higher. An example of this can be seen in Fig. 9, where for a 2ms jitter, rounding timestamps to one decimal place is optimal, while rounding has a negative impact on performance when jitter is close to zero. This also shows that, if optimal parameters are used, the impact of network jitter can be considered low in LAN scenarios, since jitter is typically below 1ms there.
@@ 5.1.3 Impact of packet loss
Packet loss presents a significant challenge to any communication protocol. Because in HCCs data is only correctly received when both the pointer and the pointed-to packet arrive at the CR, packet loss impacts protocols such as DYST and SHP more than overt traffic. In DYST, packet loss could desynchronize the CS and CR, leading to errors in message reconstruction. As discussed, SHP incorporates options for error correction tailored for the covert channel to mitigate this risk. However, high packet loss rates can still decrease effective bandwidth of the channel, as can be seen in Fig. 10. For example, a packet loss in the upper range of $10\%$, which can occur in tough WAN environments such as satellite communication [packetloss12], reduces the effective bandwidth to around $72\%$. However, typical packet loss rates, which are below $2\%$, can be considered acceptable, as the steganographic bandwidth stays over $93\%$ under these conditions.
@@ 5.1.4 Impact of overt traffic bandwidth
Given the right amount of overt traffic and a quality connection, our Python scripts achieved a fitness of up to 265bps265\;bps in live tests with amplification, reached by ISDs with $bitlength=8$ and $rehashing=7$. This far exceeds not only the performance of DYST, but also Moss, another recent network history protocol optimized for throughput [wang5634454moss]. Tab. 3 summarizes the results using these parameters under different overt traffic conditions and compares them to DYST [^1] The DYST life script was extended with the ability to record statistics and run with the parameters $chars=1,encoding=trivial$..
@@ 5.2 Detectability and Countermeasures
Established methods for detecting network timing channels generally rely on comparing packet timings with expected traffic models [detect_ipd19, NIHBook, cabuk09]. However, these approaches appear unsuitable when applied to SHP’s data channel, because it does not alter packet data or their timings. Instead, defenders must rely on detecting the signal channel that transmits the pointer.
Since our implementation of SHP uses header bits within ARP requests as the signal channel and since these bits could be modulated in arbitrary ways, the detectability evaluation will instead focus on detecting anomalies in the timing of these requests. We compare the distributions of interpacket gaps with and without SHP channels. ARP-specific detection methods, such as detecting replies without following traffic that uses the replied IP-address, are excluded here, because other protocols could be easily substituted as SHP-signaling channel (e.g., NTP [hielscher2021systematic] or HTTP [morepackets21]), but the timing of these signals would need to stay quite constant if it is to be used as a pointer.
Applicable mathematical methods for detecting timing irregularities are statistical and information-theory tests. We used the Kolmogorov-Smirnov test for the former and the compressibility score introduced by Cabuk et al. for the latter [epskappa24, keller2023improving, cabuk06]. To evaluate the detectability of SHP, we recorded real-world LAN traffic with pointers using the parameters bitlength=8, inputsource=ISD and used no extensions (out-of-order delivery, etc.).
@@ 5.2.1 KS-test
The Kolmogorov-Smirnov test is a statistical test for the similarity of two probability distributions [Iglesias2017Are, smirnov48]. It compares two empirical distribution functions $F_{n_{1}}(x)$ and $G_{n_{2}}(x)$ from two independent samples to determine if they come from the same distribution. The test statistic is the maximum absolute difference between these functions:
@@
$D$ can then be used to determine if the difference between the two distributions is statistically significant by calculating the corresponding p-value. Because this test does not assume a specific form for the underlying distributions, it can be used to compare them without assuming a particular traffic pattern [Grzesiak2023Covert, Zhang2020A].
@@ 5.2.2 Compressibility Score
The compressibility score is a statistical detection method that was first proposed for covert channels by Cabuk et al. [cabuk06, cabuk09]. This method evaluates the entropy of traffic parameters to identify patterns indicative of hidden communication. It operates on the principle that CCs often introduce information structures into otherwise random traffic, thereby reducing its overall compressibility [epskappa24]. This approach is particularly useful for identifying anomalies in sets of information that inherently exhibit high entropy, such as ARP timings.
@@
As can be expected [impactofequipment2012], for different network captures, the overt timing distributions vary widely depending on factors such as time of day, network load or equipment, and protocol behavior, while the compressibility distributions between SHP and the baseline IPDs without covert channel stayed similar to each other. This suggests that SHP packets can be introduced to the network without significantly affecting the compressibility of IPDs even under varying conditions, implying that the covert channel remains largely undetectable through this metric.
@@ 5.2.3 Detectability by Request Frequency
A different approach to detecting a CC using ARP requests is to consider the frequency of ARP requests either from a certain network node (mainly the CS) or for a certain IP or address space [Ondov2022Covert, Chen2015A]. We evaluated detectability using live SHP capture in the LAN scenario using $input=ISD$ with $bitlength=8$. Because the probability of a pointer match occurring is $1/2^{8}=1/256$, the additional request frequency introduced by the covert channel can be expected to be very low compared to the natural volume of legitimate network traffic. However, because SHP does not only use ARP as POI, but uses ARP as the signal channel, experimental verification is required. Experimental evaluations using real network traces confirm that, under the given configuration, the proportion of covert pointer requests is negligible, even if only ARP requests are considered by defenders. In several tests, the proportion of covert signals fell squarely within the typical range of legitimate traffic, making them effectively indistinguishable from normal background activity. An example can be seen in Fig. 14.
Furthermore, the stochastic nature of pointer generation under SHP means that any additional requests occur in a random pattern rather than in a regular and easily discernible rhythm. Since channels mimicking the distribution of normal traffic are significantly harder to detect [Zhang2014Network], the inherent conformity of SHP to the randomness of the network traffic makes it even less likely for statistical anomaly detection methods to isolate it based solely on request frequency [noratelimiting11]. Because SHP does not require using the requested ARP address for transmission (except for a few bits when rehashing is desired), requested addresses can be set to any desirable value to blend in with usual network activity.
@@ 5.2.5 Detectability with a GAS-style Model
To assess detectability using machine learning, we tested whether SHP’s ARP-based timing channel could be identified by a detector based on GAS, a tensor-based method for detecting covert channels that exfiltrate data through small-volume transmissions [GAS22]. The model was trained in-domain on legitimate ARP IPDs [GAS22]. This isolates the signal channel and avoids confounding domain shifts that arise when models trained on flow IPDs (e.g., HTTP/SSH) are applied to ARP.
@@
As ML-model, we adopted GAS’s sequence prediction architecture: an embedding layer over the 16 discrete states, followed by a single LSTM layer and a dense softmax output over the next-state vocabulary. We use a sequence length (time step) of 8, an embedding dimension of 100, and 64 LSTM units, consistent with [GAS22] and effective for SHP’s sparsity regime.
@@ 5.2.6 Countermeasures
One of the primary defenses against timing-based covert channels is the implementation of traffic normalization techniques [normalization20]. Protocol normalization aims to standardize network traffic characteristics to eliminate timing-based patterns. By adding jitter or enforcing fixed inter-packet intervals, the entropy of the timing information is reduced, making it more difficult for the covert sender and receiver to find matching patterns.
However, this approach introduces latency and degrades the quality of service for legitimate applications, particularly those sensitive to timing, such as VoIP or real-time streaming [Epishkina2019Timing]. For example, certain real-time applications rely on timing characteristics for synchronization or performance tuning, which normalization may disrupt [2008On]. Because of its amplification and variety of input sources, SHP can be configured to be robust while keeping the impact on bandwidth low as shown in Sect. 5.1. This means aggressive normalization would be necessary to disrupt communication. Moreover, SHP also supports input sources that are not based on IPDs such as packet order, so defenders would have to employ a wide range of normalization techniques to hamper transmission.
@@
Limiting the frequency of specific protocols, such as ARP or multicast queries, can disrupt HCCs that rely on frequent pointer signals to legitimate traffic. However, protocols like ARP exhibit traffic patterns that are inherently unpredictable due to their reactive nature, making it difficult to establish meaningful thresholds for rate limiting even traditional CC [Sun2021Suspicious, noratelimiting11]. Because SHP enables amplification of covert communication, rate limiting its communication without considerable impact on overt network usage is even less feasible.
@@
Traditional cybersecurity defense measures can also be applied to limit the effectiveness of network channels like SHP. Network segmentation can prevent broadcast or multicast communication between endpoints and minimize shared traffic accessible to HCC[Kogos2021Prospective, Ondov2022Covert]. Nevertheless, adaptive probing algorithms that analyze the communication methods available to them could pick from the parameter set of SHP, e.g., input sources and POI, to enable peer-to-peer communication over those boundaries as long as any traffic at all is still allowed [Yarochkin2009Introducing, Bortolameotti2019VictimAware].
Requiring host-based detection for all clients in a network enables detection not through the network characteristics that we have shown to be difficult to detect, but through local detection methods such as signatures or host-behaviour heuristics [Ondov2022Covert, Joe2021HostBased]. However, such measures are not any more effective against SHP than they are against other network-based CC.

### v2

@@ 5 Evaluation
Because analyzing large network captures is demanding in terms of computing resources and live experiments can only be conducted one after the other, we used genetic algorithms to find useful parameter combinations [52]. The Simple Genetic Algorithm (SGA) iteratively optimizes the selection of combinations of experimental parameters based on a performance metric computed from previous experiment runs called fitness. Each individual represents a distinct configuration of experimental parameters. At each iteration, the algorithm first selects a subset of elite individuals with the highest fitness scores. These elite configurations serve as parents in the crossover phase, where each offspring is generated by randomly inheriting parameter values from one of the two parents. To maintain diversity and explore new regions in the parameter space, a mutation operator is then applied to each offspring with a predefined mutation rate, potentially altering one or more parameter values. This straightforward evolutionary cycle—comprising selection, crossover, and mutation—enables the algorithm to progressively refine and schedule more promising parameter sets for subsequent experiments, while also ensuring that previously tested configurations are not redundantly re-evaluated.
@@
Live measurements were taken during the day and at night in a home network with 10 devices such as smartphones, smart home devices, and file servers. For offline analysis, traffic in this network was collected during the day and at night. In the WAN scenario, we used traffic recordings from the WIDE project [14].
@@ 5.1 Environmental robustness
Robustness is crucial for covert channels to ensure reliable communication in real-world network environments where such impairments are common. This is especially true for timing-based channels, as network conditions affect packet timing and order [63]. We evaluated the robustness of SHP under various network conditions, specifically focusing on the impact of delay, jitter, noise, and packet loss on bandwidth.
@@ 5.1.1 Impact of delay
The observed values for one-way network delay are typically in the ranges $\leq 110$ ms in wired LANs [42] and 5050300300 ms in WAN communication [65], depending on distance, routing, and network congestion. When using timing-based input sources, this causes differences in observed inputs between the CS and CR, especially when absolute timing is used as a reference for pointers.
In the WAN scenario, consistent network delay has a reduced impact on the robustness of SHP, as it can use relative timing (like in ISDs) or packet ordering (like in ISPNs) algorithms. This is because the CR computes intra-signal delays relative to previous pointers or initial synchronization. Let $ISD_{CS}(n)$ be the difference between when the CS observes a signal $n$ and the previous signal n1n-1, and the network delay between covert agents be $d$. Then the observed ISDs for both parties are $ISD_{CS}(n)=(t_{n}+d)-(t_{n-1}+d)=t_{n}-t_{n-1}+(d-d)=t_{n}-t_{n-1}=ISD_{CR}(n)$. This means that the observed interpacket delays or relative times remain the same, rendering any consistent network delay inconsequential for matching in the WAN scenario. This aligns with recent covert communication work in continuous or time-uncertain settings, where robustness is also achieved by leveraging timing uncertainty rather than attempting to tightly synchronize clocks [41, 36]. Similarly, for the ISPN algorithm, which relies on packet ordering, a constant delay does not alter the sequence of packet arrivals, so SHP remains robust under high network delays as long as they are stable.
@@ 5.1.2 Impact of jitter
We also implemented a silent interval using the method suggested by Wendzel et al. [62]: If the CS receives two or more POI in less than $\phi$ milliseconds, all overt traffic in that time frame will be ignored, so that only isolated POI will be considered for matching. Consequently, the CR uses $\phi$ to ignore all POI received less than this number of milliseconds before a pointer and only interpret it as pointing to POI before that. This ensures the CS can calculate and send a pointer without running into a race condition with consecutive POI. Nonetheless, the larger we choose $\phi$, the more this will reduce matching chances and thus potential bandwidth.
@@
We tested the impact of jitter using simulations as well as live network channels. In the simulations, we produced challenging network circumstances for the LAN and WAN scenarios, using a rather high time difference between CS and CR of 2ms for the LAN scenario [56] and 20ms for the WAN scenario [3] and only measured the impact of rounding network jitter of overt traffic. The results show that while less precise rounding improves performance when the jitter is negligible, more aggressive rounding improves performance when the jitter is higher. An example of this can be seen in Fig. 9, where for a 2ms jitter, rounding timestamps to one decimal place is optimal, while rounding has a negative impact on performance when jitter is close to zero. This also shows that, if optimal parameters are used, the impact of network jitter can be considered low in LAN scenarios, since jitter is typically below 1ms there.
@@ 5.1.3 Impact of packet loss
Packet loss presents a significant challenge to any communication protocol. Because in HCCs data is only correctly received when both the pointer and the pointed-to packet arrive at the CR, packet loss impacts protocols such as DYST and SHP more than overt traffic. In DYST, packet loss could desynchronize the CS and CR, leading to errors in message reconstruction. As discussed, SHP incorporates options for error correction tailored for the covert channel to mitigate this risk. However, high packet loss rates can still decrease effective bandwidth of the channel, as can be seen in Fig. 10. For example, a packet loss in the upper range of $10\%$, which can occur in tough WAN environments such as satellite communication [71], reduces the effective bandwidth to around $72\%$. However, typical packet loss rates, which are below $2\%$, can be considered acceptable, as the steganographic bandwidth stays over $93\%$ under these conditions.
@@ 5.1.4 Impact of overt traffic bandwidth
Given the right amount of overt traffic and a quality connection, our Python scripts achieved a fitness of up to 265bps265\;bps in live tests with amplification, reached by ISDs with $bitlength=8$ and $rehashing=7$. This far exceeds not only the performance of DYST, but also Moss, another recent network history protocol optimized for throughput [60]. Tab. 3 summarizes the results using these parameters under different overt traffic conditions and compares them to DYST [^1] The DYST life script was extended with the ability to record statistics and run with the parameters $chars=1,encoding=trivial$..
@@ 5.2 Detectability and Countermeasures
Established methods for detecting network timing channels generally rely on comparing packet timings with expected traffic models [40, 43, 9]. However, these approaches appear unsuitable when applied to SHP’s data channel, because it does not alter packet data or their timings. Instead, defenders must rely on detecting the signal channel that transmits the pointer.
Since our implementation of SHP uses header bits within ARP requests as the signal channel and since these bits could be modulated in arbitrary ways, the detectability evaluation will instead focus on detecting anomalies in the timing of these requests. We compare the distributions of interpacket gaps with and without SHP channels. ARP-specific detection methods, such as detecting replies without following traffic that uses the replied IP-address, are excluded here, because other protocols could be easily substituted as SHP-signaling channel (e.g., NTP [26] or HTTP [46]), but the timing of these signals would need to stay quite constant if it is to be used as a pointer.
Applicable mathematical methods for detecting timing irregularities are statistical and information-theory tests. We used the Kolmogorov-Smirnov test for the former and the compressibility score introduced by Cabuk et al. for the latter [73, 31, 10]. To evaluate the detectability of SHP, we recorded real-world LAN traffic with pointers using the parameters bitlength=8, inputsource=ISD and used no extensions (out-of-order delivery, etc.).
@@ 5.2.1 KS-test
The Kolmogorov-Smirnov test is a statistical test for the similarity of two probability distributions [27, 53]. It compares two empirical distribution functions $F_{n_{1}}(x)$ and $G_{n_{2}}(x)$ from two independent samples to determine if they come from the same distribution. The test statistic is the maximum absolute difference between these functions:
@@
$D$ can then be used to determine if the difference between the two distributions is statistically significant by calculating the corresponding p-value. Because this test does not assume a specific form for the underlying distributions, it can be used to compare them without assuming a particular traffic pattern [23, 70].
@@ 5.2.2 Compressibility Score
The compressibility score is a statistical detection method that was first proposed for covert channels by Cabuk et al. [10, 9]. This method evaluates the entropy of traffic parameters to identify patterns indicative of hidden communication. It operates on the principle that CCs often introduce information structures into otherwise random traffic, thereby reducing its overall compressibility [73]. This approach is particularly useful for identifying anomalies in sets of information that inherently exhibit high entropy, such as ARP timings.
@@
As can be expected [22], for different network captures, the overt timing distributions vary widely depending on factors such as time of day, network load or equipment, and protocol behavior, while the compressibility distributions between SHP and the baseline IPDs without covert channel stayed similar to each other. This suggests that SHP packets can be introduced to the network without significantly affecting the compressibility of IPDs even under varying conditions, implying that the covert channel remains largely undetectable through this metric.
@@ 5.2.3 Detectability by Request Frequency
A different approach to detecting a CC using ARP requests is to consider the frequency of ARP requests either from a certain network node (mainly the CS) or for a certain IP or address space [47, 13]. We evaluated detectability using live SHP capture in the LAN scenario using $input=ISD$ with $bitlength=8$. Because the probability of a pointer match occurring is $1/2^{8}=1/256$, the additional request frequency introduced by the covert channel can be expected to be very low compared to the natural volume of legitimate network traffic. However, because SHP does not only use ARP as POI, but uses ARP as the signal channel, experimental verification is required. Experimental evaluations using real network traces confirm that, under the given configuration, the proportion of covert pointer requests is negligible, even if only ARP requests are considered by defenders. In several tests, the proportion of covert signals fell squarely within the typical range of legitimate traffic, making them effectively indistinguishable from normal background activity. An example can be seen in Fig. 14.
Furthermore, the stochastic nature of pointer generation under SHP means that any additional requests occur in a random pattern rather than in a regular and easily discernible rhythm. Since channels mimicking the distribution of normal traffic are significantly harder to detect [69], the inherent conformity of SHP to the randomness of the network traffic makes it even less likely for statistical anomaly detection methods to isolate it based solely on request frequency [67]. Because SHP does not require using the requested ARP address for transmission (except for a few bits when rehashing is desired), requested addresses can be set to any desirable value to blend in with usual network activity.
@@ 5.2.5 Detectability with a GAS-style Model
To assess detectability using machine learning, we tested whether SHP’s ARP-based timing channel could be identified by a detector based on GAS, a tensor-based method for detecting covert channels that exfiltrate data through small-volume transmissions [35]. The model was trained in-domain on legitimate ARP IPDs [35]. This isolates the signal channel and avoids confounding domain shifts that arise when models trained on flow IPDs (e.g., HTTP/SSH) are applied to ARP.
@@
As ML-model, we adopted GAS’s sequence prediction architecture: an embedding layer over the 16 discrete states, followed by a single LSTM layer and a dense softmax output over the next-state vocabulary. We use a sequence length (time step) of 8, an embedding dimension of 100, and 64 LSTM units, consistent with [35] and effective for SHP’s sparsity regime.
@@ 5.2.6 Countermeasures
One of the primary defenses against timing-based covert channels is the implementation of traffic normalization techniques [54]. Protocol normalization aims to standardize network traffic characteristics to eliminate timing-based patterns. By adding jitter or enforcing fixed inter-packet intervals, the entropy of the timing information is reduced, making it more difficult for the covert sender and receiver to find matching patterns.
However, this approach introduces latency and degrades the quality of service for legitimate applications, particularly those sensitive to timing, such as VoIP or real-time streaming [18]. For example, certain real-time applications rely on timing characteristics for synchronization or performance tuning, which normalization may disrupt [4]. Because of its amplification and variety of input sources, SHP can be configured to be robust while keeping the impact on bandwidth low as shown in Sect. 5.1. This means aggressive normalization would be necessary to disrupt communication. Moreover, SHP also supports input sources that are not based on IPDs such as packet order, so defenders would have to employ a wide range of normalization techniques to hamper transmission.
@@
Limiting the frequency of specific protocols, such as ARP or multicast queries, can disrupt HCCs that rely on frequent pointer signals to legitimate traffic. However, protocols like ARP exhibit traffic patterns that are inherently unpredictable due to their reactive nature, making it difficult to establish meaningful thresholds for rate limiting even traditional CC [55, 67]. Because SHP enables amplification of covert communication, rate limiting its communication without considerable impact on overt network usage is even less feasible.
@@
Traditional cybersecurity defense measures can also be applied to limit the effectiveness of network channels like SHP. Network segmentation can prevent broadcast or multicast communication between endpoints and minimize shared traffic accessible to HCC[33, 47]. Nevertheless, adaptive probing algorithms that analyze the communication methods available to them could pick from the parameter set of SHP, e.g., input sources and POI, to enable peer-to-peer communication over those boundaries as long as any traffic at all is still allowed [66, 7].
Requiring host-based detection for all clients in a network enables detection not through the network characteristics that we have shown to be difficult to detect, but through local detection methods such as signatures or host-behaviour heuristics [47, 30]. However, such measures are not any more effective against SHP than they are against other network-based CC.